# GDSR: Global-Detail Integration through Dual-Branch Network with Wavelet Losses for Remote Sensing Image Super-Resolution

### v1

@@ Abstract
In recent years, deep neural networks, including Convolutional Neural Networks, Transformers, and State Space Models, have achieved significant progress in Remote Sensing Image (RSI) Super-Resolution (SR). However, existing SR methods typically overlook the complementary relationship between global and local dependencies. These methods either focus on capturing local information or prioritize global information, which results in models that are unable to effectively capture both global and local features simultaneously. Moreover, their computational cost becomes prohibitive when applied to large-scale RSIs. To address these challenges, we introduce the novel application of Receptance Weighted Key Value (RWKV) to RSI-SR, which captures long-range dependencies with linear complexity. To simultaneously model global and local features, we propose the Global-Detail dual-branch structure, GDSR, which performs SR reconstruction by paralleling RWKV and convolutional operations to handle large-scale RSIs. Furthermore, we introduce the Global-Detail Reconstruction Module (GDRM) as an intermediary between the two branches to bridge their complementary roles. In addition, we propose Wavelet Loss, a loss function that effectively captures high-frequency detail information in images, thereby enhancing the visual quality of SR, particularly in terms of detail reconstruction. Extensive experiments on several benchmarks, including AID, AID_CDM, RSSRD-QH, and RSSRD-QH_CDM, demonstrate that GSDR outperforms the state-of-the-art Transformer-based method HAT by an average of 0.05 dB in PSNR, while using only 63% of its parameters and 51% of its FLOPs, achieving an inference speed 2.9 times faster. Furthermore, the Wavelet Loss shows excellent generalization across various architectures, providing a novel perspective for RSI-SR enhancement.
@@
The rapid development of remote sensing technologies has resulted in an ever-increasing availability of satellite and aerial images, which play a crucial role in a variety of applications, such as environmental monitoring [1], urban planning [2], and disaster management [3]. However, the resolution of these remote sensing images is often constrained by the limitations of imaging sensors, atmospheric interference, and satellite bandwidth. Consequently, SR techniques, which seek to reconstruct high-resolution (HR) images from low-resolution (LR) observations, have garnered significant interest from both academic and industrial sectors [4, 5]. In particular, remote sensing image super-resolution (RSISR) has emerged as a pivotal research area that enhances spatial resolution and reveals finer details that are essential for accurate analysis.
Traditional SR methods have been dominated by Convolutional Neural Networks (CNNs) [6, 7] due to their strong local feature extraction and efficient computation. Models such as SRResNet [8] and EDSR [9] have demonstrated that deeper CNNs can progressively enhance SR performance. However, they suffer from the Effective Receptive Field (ERF) problem (as depicted in Fig. 1 (a)). Although deeper networks theoretically have larger receptive fields, only a small portion is effectively utilized during feature extraction, which limits their ability to capture long-range dependencies. This becomes particularly challenging in RSIs, where diverse spatial patterns and textures overwhelm the limited modeling capacity of CNN-based methods [10, 11].
In recent years, Transformers have gained significant traction in computer vision [14, 15], including RSISR [10, 16]. They excel at capturing global dependencies by modeling long-range interactions through self-attention mechanisms. However, the quadratic complexity of the self-attention operation can become a bottleneck when processing large-scale satellites or aerial images. To address this, many Transformer-based models aim to reduce computational demands by limiting attention to smaller windows [14], which in turn reduces the effective receptive field (as shown in Fig. 1 (b)) and conflicts with the goal of capturing long-range dependencies. Additionally, HAT [15] combines channel attention and window-based self-attention with an overlapping cross-attention module to capture long-range dependencies. However, although transformers can capture the global context, they struggle with computational complexity and may lose structural information during sequence transformation [17], necessitating a re-evaluation of existing methods and the development of novel approaches.
The linear architecture of Transformers like Mamba-based models [18] and RWKV-based models [13] serves as an efficient alternative to Transformers. Mamba-based models leverage State Space Models (SSMs) [19] to causally capture long-range dependencies with linear computational complexity, with recent studies suggesting that they can achieve performance comparable to or even better than Transformer-based models [18, 20]. However, their inherent sequence modeling in SSMs still poses challenges in achieving an optimal ERF for 2D images (as depicted in Fig. 1 (c)). RWKV-based models [21, 22], originally developed for natural language processing (NLP), has emerged as a potential alternative to the transformer and amba. RWKV improves efficiency with linear complexity for long-range dependencies and better captures local dependencies, enhancing performance on large-scale NLP and vision tasks. Recent studies have shown that RWKV-based models outperform Transformer-based models and linearly scaled Mamba-based models in terms of effectiveness and efficiency [22, 23]. Despite pioneering efforts to adopt RWKV for vision tasks, the adaptation of RWKV to the field of RSISR remains unexplored.
While CNNs, Transformers, Mamba, and RWKV networks each offer distinct advantages, the challenges of effectively managing receptive fields and addressing the ERF limitations remain open research questions in RSISR. Notably, a recent study [24] has shown that multi-head self-attention (MSA) and convolutional layers exhibit complementary filtering behaviors. Specifically, MSAs act as low-pass filters, emphasizing broader structures and smoothing noise, whereas conv layers serve as high-pass filters, focusing on fine-grained textures and edges. These contrasting properties make the combination of MSAs and conv layers particularly beneficial for improving model performance. Moreover, high-frequency details are crucial for accurately reconstructing remote sensing images, as they play a key role in preserving fine-grained information. However, existing approaches struggle to simultaneously model long-range dependencies while effectively capturing high-frequency details. Therefore, a natural question arises: Can a more efficient yet effective solution be developed to capture long-range dependencies across large-scale RSIs with convolution, which directly handles 2D spatial structures and offers computational advantages, while also efficiently reconstructing high-frequency details?
@@
In this paper, we propose a novel approach that integrates global representations and detailed extraction with wavelet losses to enhance the SR performance for remote sensing imagery. Our main contributions can be summarized as follows:
@@
The RWKV is introduced for remote sensing image super-resolution, along with GDSR, a novel architecture designed to effectively combine global and local information through a dual-branch structure.
@@
A Global–Detail Reconstruction Module (GDRM) is proposed to efficiently integrate information from the dual branches. By merging complementary features, this module enhances the synergy between the branches, thereby significantly improving the overall performance of RSISR tasks.
@@
Wavelet loss is introduced as a simple yet powerful loss function for RSISR. This approach effectively enhanced both the PSNR and visual quality, and demonstrated robust performance improvements across various network architectures. Moreover, its ease of implementation and adaptability make it a practical and highly effective solution for advancing RSISR techniques.
We demonstrate the empirical validity of GDSR by comparing it with various competitive CNN-, GAN-, Transformer-, Mamba-, and hybrid-based models on two RSI benchmark datasets using two degradation methods showing SOTA results. In addition, we conducted extensive ablation experiments (as detailed in Section IV-C) to investigate the approach of global-detail integration and its effect on RSI-SR, and to examine the effect of Wavelet Loss on various model architectures. We believe our work will advance efficient large-scale RSI-SR, improving both fidelity and processing speed, while ensuring robustness and practical applicability for real-world remote sensing tasks.
@@ II-A Remote Sensing Image Super-resolution
Recent advances in deep learning have greatly improved RSI-SR, with methods typically falling into CNN-based, Transformer-based, and Mamba-based categories. CNN-based methods, inspired by SRCNN [6], have used modules like residual connections [7], dense structures [25], and attention mechanisms [26], [27] to enhance performance. However, they are limited by their inability to capture long-range dependencies, which hinders their effectiveness in large-scale RSI data. Transformer-based methods, leveraging self-attention (SA) [28], have shown promise in modeling global dependencies, with approaches like window-based SA [15] and recursive SA [29] attempting to address the quadratic complexity. However, SA remains inefficient for high-resolution images, often sacrificing global modeling capability. To overcome these limitations, Mamba-based methods, which offer linear complexity for global modeling, present a compelling solution. MambaIR [18] successfully introduced Mamba into the SR domain, while FreMamba [20] pioneers the exploration of Mamba’s potential for RSI-SR, extending it with frequency analysis. Due to the unidirectional nature of SSM, Mamba’s global dependency modeling still faces challenges in achieving accurate results for RSI-SR.
@@ II-B RWKV in computer vision
Recent studies [22] indicate that RWKV can outperform both Transformers and Mamba in NLP tasks. Vision-RWKV [30] has been successfully applied to vision tasks, showing superior performance compared to vision Transformers with lower computational complexity. Several RWKV-based models, such as Diffusion-RWKV [31] for image generation and Restore-RWKV [13] for medical image restoration, have demonstrated promising results. However, the potential of RWKV for RSI-SR remains unexplored. Given that RSI exhibits diverse and complex characteristics, with texture information often subtler than in natural images, it presents unique challenges for SR tasks. Moreover, critical high-frequency details tend to be lost in deep learning models, emphasizing the need for a more effective approach to capture these crucial components. This paper pioneers the exploration of RWKV’s potential for RSI-SR by introducing a novel dual-branch architecture that combines RWKV with convolutional operations.
@@ II-C Wavelet Transforms in Deep Learning
The wavelet transform (WT) [32] has been a fundamental tool for signal processing since the 1980s and has more recently been integrated into neural network architectures for various tasks. Several studies have explored wavelet-domain methods for image restoration and SR tasks, recognizing the need to treat low- and high-frequency features differently. Methods like Wavelet-SRNet [33], WDST [34] and WDRN [35] predict wavelet coefficients to enhance SR performance. More recently, WGSR [36] introduced wavelet subband losses in GAN-based SR models to better control artifacts. In addition, WTConv [37], a novel layer leveraging the wavelet transform to effectively expand the receptive field of convolutions, has been proposed as a drop-in replacement for depth-wise convolutions in CNNs, improving performance across several key computer vision tasks. In this paper, we propose a novel Wavelet Loss that directly optimizes the reconstruction of wavelet-domain features. Our approach is designed to complement SR models based on various architectures, including CNN, GAN, Transformer, and Mamba, as well as their hybrid combinations, which have demonstrated superior performance in RSISR. Unlike previous methods, our Wavelet Loss seamlessly integrated wavelet-based supervision into the end-to-end SR training framework to ensure effective high-frequency detail recovery.
@@ III Methodology
In this section, we first introduce the overall structure of the proposed GDSR and then described three important modules in GDSR, namely RGEG, RDEG, and GDRM. Finally, we detail our designed loss function.
@@ III-A Overview of GDSR
As illustrated in Fig. 2, GDSR begins by extracting the shallow feature from the LR input using a convolutional layer. It then derives the deep features and through parallel branches: global extraction branch and details extraction branch, respectively. These features are subsequently fused using the GDRM. A global residual connection is then applied to combine the low-level and deep features, and finally, the pixel-shuffle method is employed for upsampling to generate the SR output.
@@ III-C Residual Global Extraction Group
where , , and are linear projection matrices, and is the squared ReLU activation function, which enhances nonlinearity. Notably, the transformation from to to involves a multi-layer perception (MLP) consisting of , , and , facilitating channel-wise feature fusion. The final channel mix output is obtained through:
@@ III-D Global–Detail Reconstruction Module
III-E Wavelet Loss
Due to the unique characteristics of remote sensing imagery, preserving fine-grained details across multiple scales is critical for enhancing spatial resolution while preserving fine-grained spatial details and structural consistency. To address this challenge, we introduce an integrable wavelet-based loss function designed for versatile and effective application across diverse SR models, leveraging the Haar wavelet transform to capture high-frequency details essential for visually and quantitatively improved RSI-SR results.
As illustrated in Fig. 5, the SWT generates one low-frequency (LF) subband, LL, representing the coarse structure of the image, and three high-frequency (HF) subbands, LH, HL, and HH, which encode horizontal, vertical, and diagonal details, respectively. As illustrated in Fig. 6, the wavelet loss function is formulated as the distance between the corresponding Haar wavelet subbands of the generated SR image and the HR image. The total wavelet loss is expressed as:
where and are the HR and SR images, denotes the Haar wavelet transform, indexes the subbands, and are scaling factors that balance the importance of each subband. The total wavelet loss is averaged over a minibatch size represented by . By emphasizing high-frequency subbands, the loss effectively enhances edge and texture preservation, while the low-frequency subband ensures structural consistency. The overall loss for the training with wavelet loss is given by

### v4

@@ Abstract
In recent years, deep neural networks, including Convolutional Neural Networks, Transformers, and State Space Models, have achieved significant progress in Remote Sensing Image (RSI) Super-Resolution (SR). However, existing SR methods typically overlook the complementary relationship between global and local dependencies. These methods either focus on capturing local information or prioritize global information, which results in models that are unable to effectively capture both global and local features simultaneously. Moreover, their computational cost becomes prohibitive when applied to large-scale RSIs. To address these challenges, we introduce the novel application of Receptance Weighted Key Value (RWKV) to RSI-SR, which captures long-range dependencies with linear complexity. To simultaneously model global and local features, we propose the Global-Detail dual-branch structure, GDSR, which performs SR by paralleling RWKV and convolutional operations to handle large-scale RSIs. Furthermore, we introduce the Global-Detail Reconstruction Module (GDRM) as an intermediary between the two branches to bridge their complementary roles. In addition, we propose the Dual-Group Multi-Scale Wavelet Loss, a wavelet-domain constraint mechanism via dual-group subband strategy and cross-resolution frequency alignment for enhanced reconstruction fidelity in RSI-SR. Extensive experiments under two degradation methods on several benchmarks, including AID, UCMerced, and RSSRD-QH, demonstrate that GSDR outperforms the state-of-the-art Transformer-based method HAT by an average of 0.09 dB in PSNR, while using only 63% of its parameters and 51% of its FLOPs, achieving an inference speed 3.2 times faster.
@@
The rapid development of remote sensing technologies has resulted in an ever-increasing availability of satellite and aerial images, which play a crucial role in a variety of applications, such as environmental monitoring [1], urban planning [2], and disaster management [3]. However, the resolution of these remote sensing images is often constrained by the limitations of imaging sensors, atmospheric interference, and satellite bandwidth. Consequently, SR techniques, which seek to reconstruct high-resolution (HR) images from low-resolution (LR) observations, have garnered significant interest from both academic and industrial sectors [4, 5]. In particular, remote sensing image super-resolution (RSISR) serves as a pivotal technique for enhancing spatial resolution to extract fine-grained details required for precise geospatial interpretation. However, quantitative frequency analysis highlights a fundamental imbalance in spatial frequency composition between LR and HR observations. As illustrated in Fig. 1, LR images predominantly contain low-frequency components, while critical high-frequency features representing geospatial patterns are severely attenuated. This asymmetry is exacerbated by conventional degradation processes that disproportionately suppress high-frequency spatial information, creating reconstruction bottlenecks that standard SR methods struggle to resolve effectively.
Traditional SR methods have been dominated by Convolutional Neural Networks (CNNs) [6, 7, 8, 9] due to their strong local feature extraction and computational efficiency. However, as depicted in Fig. 2 (a), the limited effective receptive field (ERF) [10] of CNNs inherently restricts their ability to model long-range dependencies. This becomes particularly challenging in large-scale RSIs, where diverse spatial patterns and textures overwhelm the limited modeling capacity of CNN-based methods [11]. While transformer-based models [12, 11, 13] alleviate the ERF limitation through global self-attention mechanisms, their quadratic computational complexity necessitates window-based self-attention that paradoxically shrinks effective receptive field coverage (as depicted in Fig. 2 (b)), and their sequence transformation processes risk losing structural information critical for spatial consistency in RSIs [14]. Recent advances in linear architectures address the computational constraints of Transformers through distinct approaches. Mamba-based models [15, 16] leverage state-space sequences (SSMs) [17] to achieve linear-complexity long-range dependency modeling, demonstrating performance comparable to Transformers. However, their inherent sequence modeling in SSMs still poses challenges in achieving an optimal ERF for 2D images (as depicted in Fig. 2 (c)), limiting spatial pattern capture in RSIs. Conversely, RWKV-based models combine linear attention mechanisms with localized dependency modeling, outperforming both Transformers and Mamba variants in vision tasks [18, 19, 20] and offering efficient long-range modeling with expanded ERF (as depicted in Fig. 2 (d)). Despite pioneering efforts to adopt RWKV for vision tasks, the adaptation of RWKV to the field of RSISR remains unexplored. Crucially, existing approaches spanning CNNs, Transformers, Mamba, and RWKV variants universally neglect the complementary synergy between global attention mechanisms (low-pass filtering) and convolutional operations (high-pass filtering) [21], failing to synergize structural context preservation with high-frequency detail recovery essential for precise large-scale RSI reconstruction. This gap underscores the need for architectures that unify efficient global dependency modeling and localized texture refinement tailored to RSISR.
@@
In this paper, we propose a novel approach that integrates global representations and detailed extraction with wavelet losses to enhance the performance of RSISR. Our main contributions can be summarized as follows:
@@
GDSR (Global-Detail Super-Resolution) is proposed to synergistically unify the RWKV and CNN within a dual-branch framework, where global contextual dependencies and localized spatial features are collaboratively refined for RSISR.
@@
A Global–Detail Reconstruction Module (GDRM) is proposed to integrate information from the dual branches efficiently. By merging complementary features, this module enhances the synergy between the branches, thereby significantly improving the overall performance of RSISR tasks.
@@
Dual-Group Multi-Scale Wavelet Loss is proposed to unify low-frequency structural subbands and aggregated high-frequency detail subbands through a dual-coefficient grouping strategy, enforcing cross-resolution frequency consistency via hierarchical wavelet decomposition across resolutions within a multi-scale framework, leveraging wavelet-domain frequency priors to augment spatial-domain optimization.
We demonstrate the empirical validity of GDSR by comparing it with various competitive CNN-, GAN-, Transformer-, Mamba-, and hybrid-based models on three RSI benchmark datasets using two degradation methods showing SOTA results. In addition, we conducted extensive ablation experiments (as detailed in Section IV-C) to investigate the approach of global-detail integration and its effect on RSI-SR, and to examine the effect of Dual-Group Multi-Scale Wavelet Loss. We believe our work will advance efficient large-scale RSI-SR, improving both fidelity and processing speed, while ensuring robustness and practical applicability for real-world remote sensing tasks.
@@ II-A Remote Sensing Image Super-resolution
Recent advances in deep learning have greatly improved RSI-SR, with methods typically falling into CNN-based, Transformer-based, and Mamba-based categories. CNN-based methods [22, 23, 24, 25], inspired by SRCNN [6], have used modules like residual connections and attention mechanisms to enhance performance. However, they are limited by their inability to capture long-range dependencies, which hinders their effectiveness in large-scale RSI data. Transformer-based methods [11, 26, 27] leverage self-attention (SA) to overcome CNN’s locality constraints. While standard SA achieves global interaction through pairwise token correlation, its quadratic complexity becomes prohibitive for high-resolution RSIs. Window-based SA [12] reduces computation via local window partitioning, yet critically limits cross-window dependencies required for reconstructing continuous geographic features (e.g., highway systems or river networks) spanning multiple windows. Mamba-based methods emerge as an alternative with linear complexity. MambaIR [15] successfully introduced Mamba [17] into the SR domain, while FreMamba [16] pioneers the exploration of Mamba’s potential for RSI-SR, extending it with frequency analysis. However, SSMs’ causal modeling framework, constrained by 1D sequence scanning of 2D spatial data, fundamentally struggles to capture omnidirectional contextual relationships required for coherent reconstruction of geometrically complex remote sensing features (e.g., roundabouts with radial roads). Emerging hybrid architectures combining multiple paradigms have demonstrated significant progress. Recent works [28, 29, 30, 31] propose innovative solutions integrating CNN-Transformer feature fusion, U-Net structures with adaptive edge recovery, and shift channel attention-enhanced Transformers, effectively addressing limitations of single-paradigm approaches while maintaining computational efficiency.
@@ II-B RWKV in computer vision
Recent studies [32, 18, 19] highlight the superiority of RWKV over Transformers and Mamba in NLP tasks due to its linear computational complexity and parallelizable recurrent architecture. In computer vision, Vision-RWKV [33] adapts RWKV for 2D spatial processing by introducing a bidirectional WKV attention mechanism to capture global dependencies and a quad-directional token shift operation to model local contextual relationships across horizontal and vertical axes. These adaptations preserve the efficiency of RWKV while enabling effective feature learning in images. Subsequent developments further expand RWKV’s applicability. Diffusion-RWKV [34] integrates RWKV into diffusion models for image generation, leveraging bidirectional processing to handle patched image inputs through linear-complexity spatial aggregation, thereby eliminating the need for window-based operations. Restore-RWKV [20] enhances 2D dependency modeling through a recurrent WKV (Re-WKV) mechanism that integrates bidirectional and recurrent attention across multiple scan directions, alongside an omnidirectional token shift (Omni-Shift) layer to broaden context aggregation.
However, existing implementations remain suboptimal for RSI due to distinct domain characteristics. Restore-RWKV’s homogeneous U-Net backbone suits anatomical medical data but falters with RSI’s extreme scale variations and heterogeneous geospatial textures. This paper introduces a dual-branch framework that decouples global feature extraction using RWKV from local feature extraction using CNN, directly addressing RSI-SR-specific challenges unmet by homogeneous architectures.
@@ II-C Wavelet Transforms in Deep Learning
The wavelet transform (WT) [35] has been a fundamental signal processing tool since the 1980s and has recently been integrated into neural network architectures. The initial wave of deep learning integration saw wavelet-based architectures like Wavelet-SRNet [36], WDST [37] and WDRN [38] employing wavelet coefficient prediction for SR enhancement. Subsequent innovations deepened this synergy, with WTConv [39] introducing trainable wavelet-based layers that dynamically expand convolutional receptive fields, demonstrating cross-task adaptability in computer vision. The integration further evolved through adversarial frameworks, particularly WGSR [40] that embedded wavelet subband losses within GAN architectures. Moreover, [41] extended this strategy by integrating wavelet and RGB losses in Transformer-based models, enhancing high-frequency preservation and achieving superior visual performance on natural images compared to pixel-wise loss training.
Notably, existing wavelet-based SR methods typically process subbands at a single scale, failing to capture the multi-scale characteristics intrinsic to RSI textures. This fragmented optimization approach separately constrains low-frequency structural components and high-frequency directional details without explicit cross-scale coordination, leading to inconsistent reconstruction of hierarchical spatial patterns. In this paper, we propose a novel Dual-Group Multi-Scale Wavelet Loss to enforce cross-resolution frequency consistency through explicit approximation subband grouping and multi-scale high-frequency aggregation (horizontal/vertical/diagonal).
@@ III Methodology
In this section, we first introduce the overall structure of the proposed GDSR and then describe three important modules in GDSR, namely RGEG, RDEG, and GDRM. Finally, we detail our designed loss function.
@@ III-A Overview of GDSR
As illustrated in Fig. 3, GDSR begins by extracting the shallow feature from the LR input using a convolutional layer. It then derives the deep features and through parallel branches: global extraction branch and details extraction branch, respectively. These features are subsequently fused using the GDRM. A global residual connection is then applied to combine the low-level and deep features, and finally, the pixel-shuffle method [42] is employed for upsampling to generate the SR output.
@@ III-C Residual Global Extraction Group
where , , and are linear projection matrices, and is the squared ReLU activation function, which enhances nonlinearity. Notably, the transformation from to to involves a multi-layer perception consisting of , , and , facilitating channel-wise feature fusion. The final channel mix output is obtained through:
@@ III-D Global–Detail Reconstruction Module
III-E Dual-Group Multi-Scale Wavelet Loss
The proposed wavelet loss construction follows the paradigm established in prior studies [40, 41]. As illustrated in Fig. 6, initially, both the HR image and SR image are transformed into the YCbCr color space. The stationary wavelet transform (SWT) is subsequently applied exclusively to the luminance (Y) channel, as human visual perception demonstrates higher sensitivity to structural details in this component. Through SWT decomposition, one low-frequency (LF) subband (LL) and three high-frequency (HF) subbands (LH, HL, HH) are generated. The LL subband encapsulates coarse structural information, while the HF components encode directional details: LH captures horizontal features, HL preserves vertical characteristics, and HH retains diagonal patterns. These HF subbands are concatenated along the channel dimension to form a composite HF representation , whereas the LL subband is retained as .
The wavelet loss function is formulated as the distance between the corresponding subbands of the generated SR image and the HR image. The total wavelet loss is expressed as:
where denotes three spatial resolutions, , and , represent the LF/HF subbands of the SR/HR images at scale , are resolution-specific weights satisfying (Specific settings for can be found in Table VI, but generally, arbitrary can be used), and , control the global contributions of LF/HF components. The overall loss for the training with wavelet loss is given by
@@ IV Experiment
IV-A Datasets and Evaluation
This paper reports the performance of SR on three RSI datasets, including the publicly available AID [44], UCMerced [45], and the custom RSSRD-QH dataset.
The AID dataset consists of 30 categories, with the data in each category randomly divided into two equal parts for training and testing. Additionally, 20% of the training data was reserved as a validation set, resulting in 4,000 training images, 1,000 validation images, and 5,000 test images. Similarly, the UCMerced dataset, comprising 21 categories with a total of 2,100 images, was processed identically, yielding 945 training images, 105 validation images (10% of the training data), and 1,050 test images.
The RSSRD-QH dataset is located in representative watershed units within the Sanjiangyuan region and the Qinghai Lake surrounding area. The geographical distribution of the counties covered by the dataset, including their longitude and latitude ranges, is summarized in Table I. The dataset was collected using unmanned aerial vehicles, with flight altitudes carefully controlled between 30 and 50 meters. The average spatial resolution of the captured imagery is 0.01 meters, providing comprehensive coverage of various landscape elements across different elevation gradients, including terraces, floodplains, gentle slopes, and steep slopes within these watershed units. From each Watershed, we randomly selected 50% of the images for the training set and 50% for the test set, with 20% of the training set randomly selected for validation. This resulted in a training set of 2,242 images, a validation set of 561 images, and a test set of 2,804 images. The HR images in the AID and RSSRD-QH datasets were 600 × 600 pixels, while those in the UCMerced dataset were 256 × 256 pixels.
The proposed method, along with other competing methods, was evaluated on the test datasets using two classic full-reference metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [46]. Additionally, the LPIPS [47] metric was used to capture the perceptual quality of the images. Notably, PSNR and SSIM were calculated based on the luminance channel (Y) of the YCbCr color space, while LPIPS evaluated the perceptual similarity between images by extracting features from a pre-trained deep network, specifically the AlexNet [48], which was employed in this study.
IV-B Degradation Model and Implementation Details
We employed both bicubic downsampling and a comprehensive degradation model (CDM) to simulate LR images during training. Based on existing blind SR methods [49, 50] and the characteristics of remote sensing sensors [51, 52], the CDM for LR image synthesis included isotropic Gaussian blur, motion blur, scaling with different interpolation methods, additive Gaussian noise, and JPEG compression noise, all applied in random order.
Our GDSR performed deep feature exploration through 4 RGEGs, RDEGs, and GDRMs. Each RGEG consisted of 6 R-RWKVBs, and each RDEG contains 12 RCBs. Empirically, we set the internal channel dimension to = 96. For the GDSR_TC and GDSR_MC, the RGEG branches correspond to the RTSB and ResidualGroup of SwinIR [12] and MambaIR [15], respectively, with all other settings remaining consistent with GDSR.
During training, 64 × 64 patches were randomly cropped from LR RSIs, along with their HR counterparts scaled by the target scale factor (×2, ×3, ×4). The batch size was set to 16. The initial learning rate was set to and optimized using the Adam optimizer, where = 0.9 and = 0.99, with a step learning rate scheduler halving the rate at the 100th epoch over a total of 200 training epochs. All SR models were implemented in the PyTorch framework. The FLOPs results were calculated using an input tensor of size 1 × 3 × 160 × 160, and the computations were performed using torch-operation-counter 111https://github.com/SamirMoustafa/torch-operation-counter. The inference times were tested on 300 random images with a size of 3 × 160 × 160. All experiments were conducted on an NVIDIA A800 80GB GPU.